Prove that the following statements are true by means of mathematical induction n X


(a) ∀n ≥ 1, i=1 i(i!) = (n + 1)! − 1.


(b) 7n −2n is divisible by 5, ∀n ≥ 0.


(c) n! > n2, ∀n ≥ 4.

1
Expert's answer
2021-11-05T00:33:45-0400

a)

"\displaystyle{\sum^n_{i=1} }i(i!)=(n-1)!-1"